OpenStudy (anonymous):

Whats the sum of the first 50 terms? \( \huge { 12 - -\frac{7}{10}n } \) I am getting \ ( \huge -\frac{445}{2} \) is that right because it is stating it is wrong. They stat it is \( \huge -\frac{585}{2} \) I am using \( \huge \frac{n}{2}[2*a_1 + (n-1)d) ] to solve\)

OpenStudy (anonymous):

Whats the sum of the first 50 terms? \( \huge { 12 - -\frac{7}{10}n } \) I am getting \ ( \huge -\frac{445}{2} \) is that right because it is stating it is wrong. They stat it is \( \huge -\frac{585}{2} \) I am using \( \huge \frac{n}{2}[2*a_1 + (n-1)d) ] \) to solve

OpenStudy (anonymous):

\( \huge \frac{50}{2}[2*\frac{127}{10} + (50-1)*-\frac{7}{10} ]\)

OpenStudy (anonymous):

This \( \huge \huge { 12 - -\frac{7}{10}n }\) SHOULD BE \( \huge\huge { 12 - \frac{7}{10}n } \)

OpenStudy (anonymous):

Wait, so is the sequence \(a_n=12-\frac{7}{10}n\), \(n>0\)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok, if so, I think what you did wrong is that your first term is wrong.\[a_1=12-\frac{7}{10}=\frac{113}{10}\]

OpenStudy (anonymous):

Ah and I got 127/10. Now how did I get 127...... That was the problem.... I hate mis calculating problems lol Thank you. It came up to be correct.

OpenStudy (anonymous):

lol np. Everyone makes these kinds of mistakes from time to time. :)